There were 5 times as many boys as girls at the party. Also, the number of boys was 100 less than 15 times the number of girls. How many boys and girls came to the party?
Answer by LilSkittleMd(119)

Let b=number of boys
g=number of girls
The equations are:
b=5g
b=100-15g
Since both equations are equal to b,they are equal to each other:
5g=100-15g
Now solve the equation
20g=100
g=5
Substitue the number of girls in one of the original equations
b=5(5)
b=25
There were 5 girls at the party and 25 boys at the party